### v2

@@ 4 Consistent conjectures
It is natural to wonder how permissive consistency is as a requirement. We now illustrate via 4.1 that the conjectures in the rational expectations solution do not satisfy consistency.Footnote 11Footnote 11Footnote 11Because consistency is only a sufficient condition, this is not enough to conclude the rational expectations solution may be empty. In recent work, Nicolò et al. (2024) show the rational expectations solution may be empty in dynamic matching markets. Before introducing the example, it is worth defining the rational expectations solution in the context of the dynamic economy. To that end, fix an economy $E^{T}$ and a period-1 agent, $k\in A_{1}\cup B_{1}$. Define the length-$T$ economy, $E^{T}_{\setminus k}$ to be an economy that coincides with $E^{T}$ except that agent $k$ arrives in period $2$.Footnote 12Footnote 12Footnote 12Formally, if $k\in A_{1}$, $E^{T}_{\setminus k}=(A_{1}\setminus\{k\},B_{1},A_{2}\cup\{k\},B_{2},\dots,A_{T},B_{T})$. Rational expectations presumes we have already defined the rational expectations solutions for $E^{T}_{\setminus k}$, denote this by $\mathrm{RE}_{T}(E^{T}_{\setminus k})$. Under rational expectations, agent $k$ believes that when they decide to be unavailable to match in period $1$, the matching that ensues in the economy from period $1$ onward is governed by the solution concept while respecting $k$’s decision to not participate in the period-1 matching. Formally, agent $k$’s conjectures are given by $\mathrm{RE}_{T}(E^{T}_{\setminus k})$.Footnote 13Footnote 13Footnote 13Note the slight abuse of notation. We defined conjectures in 5 as a subset of $\pazocal{M}_{T}(\emptyset,k)$ so that a conjecture specifies that agent $k$ is unmatched in period $1$. Instead, $\mathrm{RE}_{T}(E^{T}_{\setminus k})$ is a subset of the set of matchings for $E^{T}_{\setminus k}$ in which $k$ only arrives in period $2$ and hence, specifying $k$’s outcome in period $1$ is not necessary.
@@
Arrivals are given by $A_{1}=\{a_{1},a_{2},a_{3}\},B_{1}=\{b_{1},b_{2}\},A_{2}=\{a_{4}\},B_{2}=\{b_{3},b_{4}\}$. Below, we list the agents’ preferences. If $(b,0)$ appears before $(b^{\prime},1)$ in the ranking of an agent on side $A$, then they prefer to match immediately with $b$ than wait one period to match with $b^{\prime}$. That is, the 00s and $1$s are the exponents of the discount factors, and the list provides the ranking of the discounted utilities $\{\delta_{\cdot}^{t-1}u(\cdot,b):a\in A_{1}\}$. For the $t=1$-agents for whom we present the ranking of utilities alone (i.e., $a_{2},b_{1}$), we assume the discount factor is close to $1$–though not exactly $1$–so that the (static) rankings also represent the rankings over the dynamic matchings.
@@
Furthermore, $m^{\star}$ is a rational expectations solution for the economy, which is supported by the following conjectures.Footnote 14Footnote 14Footnote 14We provide the supporting details for this example in Appendix A. Agent $a_{2}$ cannot block $m^{\star}$ because
@@
However, the conjectures do not satisfy consistency because $m^{\star}$ is not a valid conjecture for $a_{3}$: In every rational expectations solution in $E_{\setminus\{a_{3},b_{1}\}}^{2}$, $b_{1}$ is matched with $a_{4}$, making it impossible to support a solution in $E_{\setminus\{a_{3}\}}^{2}$ which has $b_{1}$ matched with $a_{1}$. Because $m^{\star}$ is a rational expectations solution for the economy, it follows that the solution concept also fails generalized consistency.
The failure of consistency in the rational expectations solution can be tied back to the timing assumptions implicit in the solution concept and the conjectures. Given their conjectures, the solution concept presumes the $t=1$-agents decide simultaneously whether to be available to match in $t=1$. Instead, each agent’s conjectures presume they are the first to decide not to be available to match in $t=1$, while the remaining contemporaneous agents remain available to match in $t=1$. For instance, when $b_{1}$ considers not being available to match in $t=1$, $b_{1}$ presumes $a_{3}$ is available to match in $t=1$–and in fact, $a_{3}$ must match in $t=1$ to prevent $b_{1}$ from blocking $m^{\star}$. Instead, $a_{3}$’s conjectures presume $a_{3}$ cannot match in $t=1$, and given this, $b_{1}$ is able to match with $a_{4}$ in $t=2$. For this reason, $m^{\star}$ cannot be compatible with $a_{3}$’s conjectures: supporting $m^{\star}$ as a rational expectations solution requires that $a_{3}$ may possibly match in $t=1$.
The analysis so far identified a property of the family of conjectures, $\Phi$, that guarantees existence, implicitly separating what the solution concept describes “on path” from the conjectures that support this path. Instead, we have said little about what these conjectures entail. For instance, whereas consistency implies the conjectures are consistent with the solution concept “on the path” (i.e., for the agents who remain unmatched), it does not require that the same holds “off the path” (i.e., for the agents who are matched). Consequently, agents could be ascribing to $m^{\star}$ for fear of facing continuation matchings that are, in some sense, not credible; a property that would make the $\Phi$-solutions less appealing. Conjectures such as those in Equation 4.1 satisfy consistency and avoid these “non-credible” threats. However, they impose no structure on the matching among the period-1 agents. Section 5 introduce two families of conjectures (and their corresponding solution concepts) addressing these issues.
@@ 5 Two solution concepts with consistent conjectures
In this section, we introduce two families of consistent conjectures and their corresponding solution concepts. Section 5.1 defines CVR-dynamic stability, a refinement of dynamic stability in Doval (2022). Section 5.2 defines sophisticated dynamic stability, an extension to dynamic matching markets of the eponymous solution concept in Hafalir (2008).
@@ 5.1 Continuation-values-respecting dynamic stability
The solution concept we introduce in this section, CVR-dynamic stability, builds on the conjectures in Equation 4.1. Whereas the conjectures in Equation 4.1 require agents to believe that the continuation matching abides by the solution concept, they place no restriction on what period-$t$ matching an agent expects would ensue conditional on not matching in period $t$. CVR-dynamic stability imposes a minimal requirement on the period-$t$ matching that reflects other agents’ continuation values. In what follows, we first define the requirement on the period-$t$ matching (9), the conjectures for CVR-dynamic stability (see Equation 5.3), and the solution concept (10). We then show in 5.1 that CVR-dynamic stability has consistent conjectures. We end the section by discussing the connection with dynamic stability in Doval (2022).
To define CVR-dynamic stability, we first introduce a condition on the period-$t$ matching. To that end, fix an economy, $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Suppose that for each agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we are given a set of matchings $\varphi_{k}\subset\pazocal{M}_{T}(m^{t-1},k)$. Similar to 8, we can define the one-period economy with agents in $\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ induced by $\{\varphi_{k}:k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$, which we denote by $E_{t,\varphi}$. Given a period-$t$ matching, $m_{t}$, let $E_{t,\varphi}^{m_{t}}$ denote the restriction of $E_{t,\varphi}$ to the set of agents $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m_{t}(k)\neq k$.
@@
The period-$t$ matching, $m_{t}$, is stable among those who match in period $t$ relative to $\varphi$ if $m_{t}\in\mathrm{S}(E_{t,\varphi}^{m_{t}})$.
@@
As the name suggests, 9 requires the period-$t$ matching satisfies a condition analogous to the static notion of stability among the agents who exit the economy in period $t$. We say analogous because rather than requiring that the period-$t$ matching be individually rational, 9 requires each agent $k$ be matched to a partner they prefer to the worst matching in $\varphi_{k}$.Footnote 15Footnote 15Footnote 15 If for each agent $k$, we take $\varphi_{k}$ to be the subset of individually rational matchings for $k$ in $\pazocal{M}_{T}(m^{t-1},k)$, we obtain the condition in Doval (2022, Definition 4). We discuss this in detail below.
@@
Given 9, we now define the conjectures associated with CVR-dynamic stability. Suppose we have already defined CVR-dynamic stability for economies of length $\{1,\dots,T-t\}$ and let $\hat{\mathrm{DS}}_{T-t}$ denote the correspondence of $\mathrm{CVR}$-dynamically stable matchings for length $T-t$ economies. Fix agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ and define their conjectures under CVR-dynamic stability as follows:
@@
$\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):\begin{array}[]{ll}\text{ (i) }\overline{m}_{t}\in\mathrm{S}\left(E_{t,\varphi_{\hat{\mathrm{DS}},t}}^{\overline{m}_{t}}\right)\\ \text{ (ii) }(\overline{m}_{s})_{s=t+1}^{T}\in\hat{\mathrm{DS}}_{T-t}(E^{T}_{t+1}(\overline{m}^{t}))\end{array}\right\}.$ (5.3)
@@
That is, when considering not being available to match in period $t$, $k$’s conjectured matchings satisfy two properties. First, like the conjectures in Equation 4.1, $k$ assumes that the continuation matching will be a $\mathrm{CVR}$-dynamically stable matching. Second, $k$ conjectures that those agents who exit the economy in period $t$ under $\overline{m}_{t}$ could not have found a better matching among themselves. That is, $\overline{m}_{t}$ cannot be improved by pairwise blocks, and no agent who matches in period $t$ matches with someone worse than their implicit continuation value according to their own conjectures, $\varphi_{\hat{\mathrm{DS}},t}$.Footnote 16Footnote 16Footnote 16The last part is consistent with the idea that agents simultaneously decide whether to match or remain unmatched given their set of conjectures (recall the discussion in 4.1). This last aspect of $k$’s conjectures is the sense in which they respect other contemporaneous agents’ continuation values (as defined by their conjectures). Despite the fixed-point flavor of the set in Equation 5.3, agent $k$’s conjectures are well-defined: Because $k$ is unmatched in period $t$ under $\overline{m}$, the set $\mathrm{S}\left(E_{t,\varphi_{\hat{\mathrm{DS}},t}}^{\overline{m}_{t}}\right)$ can be defined independently of agent $k$’s conjectures.
@@
We can define the set of $\mathrm{CVR}$-dynamically stable matchings for $E^{T}_{t}(m^{t-1})$, $\hat{\mathrm{DS}}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},\cdot)$ as the conjectures and $\hat{\mathrm{DS}}_{T-t}$ as the continuation solution concept. Working recursively in this way, we can define the family of conjectures $\Phi_{\hat{\mathrm{DS}}}$ for $E^{T}$ to be the collection $\{\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k):t\geq 1,m^{t-1},k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$. The definition of CVR-dynamic stability is now immediate:
Definition 10 (CVR-dynamic stability).
@@
A matching $m$ for $E^{T}$ is $\mathrm{CVR}$-dynamically stable if it is a $\Phi_{\hat{\mathrm{DS}}}$-solution for $E^{T}$. We denote by $\hat{\mathrm{DS}}_{T}(E^{T})$ the set of such solutions.
@@
The proof of 5.1 is in Appendix B. The proof proceeds by induction on the length of the economy, first showing that the set of conjectures is nonempty and then showing they satisfy consistency; Theorem 4.1 then implies the set of $\mathrm{CVR}$-dynamically stable matchings is nonempty. Thus, the key step is showing that the conjectures are nonempty: Because of the fixed-point character of the conjectures in period $t$, we build a recursion whose iteration delivers the set $\varphi_{\hat{\mathrm{DS}},t}$ as a limit.
@@
CVR-dynamic stability is a refinement of dynamic stability, the stability notion in Doval (2022). In particular, if we denote by $\varphi_{\mathrm{DS}}$ the conjectures corresponding to dynamic stability, one can show that for all $t$, all $m^{t-1}$, and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, the conjectures $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k)\subseteq\varphi_{\mathrm{DS},t}(m^{t-1},k)$, which in turn implies the set of $\mathrm{CVR}$-dynamically stable matchings is a subset of those that are dynamically stable.
To understand the difference between the solution concepts, it is instructive to compare the conjectures in Equation 5.3 to those in Doval (2022). Whereas both solution concepts require agents to conjecture that the continuation matching conforms with the solution concept in the induced continuation economy (i.e., requirement (ii) in Equation 5.3), the conjectures in Doval (2022) impose a weaker requirement on the period-$t$ matching. Namely, that among the agents who exit the economy in period $t$, the period-$t$ matching is individually rational and no blocking pairs exist.Footnote 17Footnote 17Footnote 17Formally, let $\varphi_{k}^{IR}$ denote the set of matchings in $\pazocal{M}_{T}(m^{t-1},k)$ that are individually rational for $k$. Then, in dynamic stability, matching $\overline{m}\in\pazocal{M}_{T}(m^{t-1},k)$ is a conjecture for $k$ if (i) $\overline{m}_{t}$ is stable among those who match relative to continuation values $\varphi_{\cdot}^{IR}$ and (ii) the continuation matching is dynamically stable. In other words, under dynamic stability, an agent $k$ conjectures that other contemporaneous agents are willing to exit in period $t$ with partners that dominate being unmatched through period $T$, even if they may not dominate the worst conjectured matching under $\varphi_{\mathrm{DS},t}(m^{t-1},\cdot)$. At the same time, when considering whether to block the proposed matching in period $t$, the agents available to match in period $t$ compare their payoff under the proposed matching to their worst conjectured matching under $\varphi_{\mathrm{DS},t}(m^{t-1},\cdot)$. That is, in dynamic stability, agents make their decisions to be available to match in period $t$ relative to their conjectures, but these conjectures do not take into account that other contemporaneous agents use their own conjectures when evaluating whether to match in period $t$. CVR-dynamic stability eliminates this gap by requiring the period-$t$ agents to internalize that other contemporaneous agents’ continuation values are exactly those prescribed by their conjectures.
We conclude this section illustrating that CVR-dynamic stability is a strict refinement of dynamic stability:
@@
The matching on the left, $m^{L}$, is a dynamically stable matching, with $m^{C}$ describing the conjectured matching that dissuades $a_{2}$ from waiting to be matched in $t=2$. Importantly, under $m^{C}$, $a_{1}$ and $b_{1}$ match in period $1$. However, it is easy to see that $a_{1}$ can always guarantee matching with $b_{3}$ (and similarly, $b_{1}$ with $a_{3}$) by remaining unmatched in $t=1$. In other words, $m^{C}$ cannot be a conjecture for $a_{2}$ under $\varphi_{\hat{\mathrm{DS}}}$ and hence, $m^{L}$ is not $\mathrm{CVR}$-dynamically stable. Instead, $m^{R}$ is $\mathrm{CVR}$-dynamically stable (and hence, dynamically stable).

### v3

@@ 4 Consistent conjectures
A natural question is how permissive consistency is as a requirement. We illustrate via 4.1 that the conjectures in the rational expectations solution do not satisfy consistency.Footnote 11Footnote 11Footnote 11Because consistency is only a sufficient condition, that the conjectures in the rational expectations solution fail consistency is not enough to conclude the rational expectations solution may be empty. In recent work, Nicolò et al. (2024) show the rational expectations solution may be empty in dynamic matching markets. Before introducing the example, defining the rational expectations solution in the context of the dynamic economy is useful; we do so informally here and formally in Section A.1. To that end, fix an economy $E^{T}$ and a period-1 agent, $k\in A_{1}\cup B_{1}$. Define the length-$T$ economy, $E^{T}_{\setminus k}$, to be an economy that coincides with $E^{T}$ except that agent $k$ arrives in period $2$.Footnote 12Footnote 12Footnote 12Formally, if $k\in A_{1}$, $E^{T}_{\setminus k}=(A_{1}\setminus\{k\},B_{1},A_{2}\cup\{k\},B_{2},\dots,A_{T},B_{T})$. The definition of the rational expectations solution for $E^{T}$  builds on the rational expectations solution for $E^{T}_{\setminus k}$, which we denote by $\mathrm{RE}_{T}(E^{T}_{\setminus k})$. Under rational expectations, when agent $k$ considers not being available to match in period $1$, they anticipate that the matching that ensues from period $1$ onward is governed by the solution concept, while respecting $k$’s decision to not participate in the period-1 matching. Formally, agent $k$’s conjectures are given by $\mathrm{RE}_{T}(E^{T}_{\setminus k})$.Footnote 13Footnote 13Footnote 13Note the slight abuse of notation. We defined conjectures in 5 as a subset of $\pazocal{M}_{T}(\emptyset,k)$ so that a conjecture specifies that agent $k$ is unmatched in period $1$. Instead, $\mathrm{RE}_{T}(E^{T}_{\setminus k})$ is a subset of the set of matchings for $E^{T}_{\setminus k}$ in which $k$ only arrives in period $2$ and hence, specifying $k$’s outcome in period $1$ is not necessary. The rational expectations solution for $E^{T}$ is the set of matchings that satisfy 7 when for all $k\in A_{1}\cup B_{1}$, conjectures are determined by $\mathrm{RE}_{T}(E^{T}_{\setminus k})$ and continuation matchings are elements of $\mathrm{RE}_{T-1}(\cdot)$.
@@
Arrivals are given by $A_{1}=\{a_{1},a_{2},a_{3}\},B_{1}=\{b_{1},b_{2}\},A_{2}=\{a_{4}\},B_{2}=\{b_{3},b_{4}\}$. Below, we list the agents’ preferences. If $(b,0)$ appears before $(b^{\prime},1)$ in the ranking of an agent on side $A$, they prefer to match immediately with $b$ than wait one period to match with $b^{\prime}$. That is, the 00s and $1$s are the exponents of the discount factors, and the list provides the ranking of the discounted utilities $\{\delta_{\cdot}^{t-1}u(\cdot,b):a\in A_{1}\}$. For the $t=1$-agents for whom we present the ranking of utilities alone (i.e., $a_{2},b_{1}$), we assume the discount factor is close to $1$—though not exactly $1$—so that the (static) rankings also represent the rankings over the dynamic matchings.
@@
Furthermore, $m^{\star}$ is a rational expectations solution for the economy, which is supported by the following conjectures.Footnote 14Footnote 14Footnote 14We provide the supporting details for this example in Section A.1. Agent $a_{2}$ cannot block $m^{\star}$ because
@@
However, the conjectures do not satisfy consistency, because $m^{\star}$ is not a valid conjecture for $a_{3}$. In every rational expectations solution in $E_{\setminus\{a_{3},b_{1}\}}^{2}$, $b_{1}$ is matched with $a_{4}$, making supporting a solution in $E_{\setminus\{a_{3}\}}^{2}$, which has $b_{1}$ matched with $a_{1}$, impossible. Because $m^{\star}$ is a rational expectations solution for the economy, it follows that the solution concept also fails generalized consistency.
The failure of consistency in the rational expectations solution can be tied to the timing assumptions implicit in the solution concept and the conjectures. Given their conjectures, the solution concept presumes the $t=1$-agents decide simultaneously whether to be available to match in $t=1$. Instead, each agent’s conjectures presume they are the first to decide not to be available to match in $t=1$, whereas the remaining contemporaneous agents remain available to match in $t=1$. For instance, when $b_{1}$ considers not being available to match in $t=1$, $b_{1}$ presumes $a_{3}$ is available to match in $t=1$, and in fact, $a_{3}$ must match in $t=1$ to prevent $b_{1}$ from blocking $m^{\star}$. Instead, $a_{3}$’s conjectures presume $a_{3}$ cannot match in $t=1$, and thus, $b_{1}$ is able to match with $a_{4}$ in $t=2$. For this reason, $m^{\star}$ cannot be compatible with $a_{3}$’s conjectures: supporting $m^{\star}$ as a rational expectations solution requires that $a_{3}$ may possibly match in $t=1$.
The analysis has so far identified a property of the family of conjectures, $\Phi$, that guarantees existence. Yet, we have said little about what these conjectures entail. For instance, the conjectures could include continuation matchings that are, in some sense, not credible; a property that would make the $\Phi$-solutions less appealing. Conjectures such as those in Equation 4.1 satisfy consistency and avoid these “non-credible” threats. However, they impose no structure on the matching among the period-1 agents. Section 5 introduces two families of conjectures (and their corresponding solution concepts) addressing these issues.
@@ 5 Two solution concepts with consistent conjectures
In this section, we introduce two families of consistent conjectures and their corresponding solution concepts. Section 5.1 defines $\mathrm{CVR}$ dynamic stability, a refinement of dynamic stability in Doval (2022). Section 5.2 defines sophisticated dynamic stability, an extension to dynamic matching markets of the eponymous solution concept in Hafalir (2008).
@@ 5.1 Continuation-values-respecting dynamic stability
The solution concept we introduce in this section, $\mathrm{CVR}$ dynamic stability, builds on the conjectures in Equation 4.1. Whereas the conjectures in Equation 4.1 require agents to believe the continuation matching abides by the solution concept, they place no restriction on what period-$t$ matching an agent expects would ensue conditional on not matching in period $t$. $\mathrm{CVR}$ dynamic stability imposes a minimal requirement on the period-$t$ matching that reflects other agents’ continuation values. In what follows, we first define the requirement on the period-$t$ matching (10), the conjectures for $\mathrm{CVR}$ dynamic stability (see Equation 5.3), and the solution concept (11). We then show in 5.1 that $\mathrm{CVR}$ dynamic stability has consistent conjectures. We end the section by discussing the connection with dynamic stability in Doval (2022).
To define $\mathrm{CVR}$ dynamic stability, we first introduce a condition on the period-$t$ matching. To that end, fix an economy, $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Suppose that for each agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we are given a set of matchings $\varphi_{k}\subset\pazocal{M}_{T}(m^{t-1},k)$. Similar to 8, we can define the one-period economy with agents in $\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ induced by $\{\varphi_{k}:k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$, which we denote by $E_{t,\varphi}$. Given a period-$t$ matching, $m_{t}$, let $E_{t,\varphi}^{m_{t}}$ denote the restriction of $E_{t,\varphi}$ to the set of agents $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m_{t}(k)\neq k$.
@@
The period-$t$ matching, $m_{t}$, is stable among those who match in period $t$ relative to $\varphi$ if the restriction of $m_{t}$ to $E_{t,\varphi}^{m_{t}}$ is an element of $\mathrm{S}(E_{t,\varphi}^{m_{t}})$, that is, $m_{t}|_{E_{t,\varphi}^{m_{t}}}\in\mathrm{S}(E_{t,\varphi}^{m_{t}})$.
@@
As the name suggests, 10 requires that the period-$t$ matching satisfies a condition analogous to the static notion of stability among the agents who exit the economy in period $t$. We say “analogous” because rather than requiring that the period-$t$ matching be individually rational, 10 requires that each agent $k$ be matched to a partner they prefer to the worst matching in $\varphi_{k}$.Footnote 15Footnote 15Footnote 15 If for each agent $k$, we take $\varphi_{k}^{IR}$ to be the subset of individually rational matchings for $k$ in $\pazocal{M}_{T}(m^{t-1},k)$, we obtain the condition in Doval (2022, Definition 4). We discuss the connection with Doval (2022) in detail below.
@@
Given 10, we now define the conjectures associated with $\mathrm{CVR}$ dynamic stability. Suppose we have already defined $\mathrm{CVR}$ dynamic stability for economies of length $\{1,\dots,T-t\}$ and let $\hat{\mathrm{DS}}_{T-t}$ denote the correspondence of $\mathrm{CVR}$ dynamically stable matchings for length $T-t$ economies. Fix agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ and define their conjectures under $\mathrm{CVR}$ dynamic stability as follows:
@@
$\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):\begin{array}[]{ll}\text{ (i) }\overline{m}_{t}|_{E_{t,\varphi_{\hat{\mathrm{DS}},t}}^{\overline{m}_{t}}}\in\mathrm{S}\left(E_{t,\varphi_{\hat{\mathrm{DS}},t}}^{\overline{m}_{t}}\right)\\ \text{ (ii) }(\overline{m}_{s})_{s=t+1}^{T}|_{E^{T}_{t+1}(\overline{m}^{t})}\in\hat{\mathrm{DS}}_{T-t}(E^{T}_{t+1}(\overline{m}^{t}))\end{array}\right\}.$ (5.3)
@@
That is, when considering not being available to match in period $t$, $k$’s conjectured matchings satisfy two properties. First, $k$ assumes the continuation matching will be a $\mathrm{CVR}$ dynamically stable matching. Second, $k$ conjectures that those agents who exit the economy in period $t$ under $\overline{m}_{t}$ could not have found a better matching among themselves. That is, $\overline{m}_{t}$ cannot be improved by pairwise blocks, and no agent who matches in period $t$ matches with someone worse than their implicit continuation value according to their own conjectures, $\varphi_{\hat{\mathrm{DS}},t}$.Footnote 16Footnote 16Footnote 16The last part is consistent with the idea that agents simultaneously decide whether to match or remain unmatched given their set of conjectures (recall the discussion in 4.1). This last aspect of $k$’s conjectures is the sense in which they respect other contemporaneous agents’ continuation values (as defined by their conjectures). Despite the fixed-point flavor of the set in Equation 5.3, agent $k$’s conjectures are well defined: because $k$ is unmatched in period $t$ under $\overline{m}$, the set $\mathrm{S}\left(E_{t,\varphi_{\hat{\mathrm{DS}},t}}^{\overline{m}_{t}}\right)$ can be defined independently of agent $k$’s conjectures.
@@
We can define the set of $\mathrm{CVR}$ dynamically stable matchings for $E^{T}_{t}(m^{t-1})$, $\hat{\mathrm{DS}}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},\cdot)$ as the conjectures and $\hat{\mathrm{DS}}_{T-t}$ as the continuation solution concept. Working recursively in this way, we can define the family of conjectures $\Phi_{\hat{\mathrm{DS}}}$ for $E^{T}$ to be the collection $\{\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k):t\geq 1,m^{t-1},k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$. The definition of $\mathrm{CVR}$ dynamic stability is now immediate:
Definition 11 ($\mathrm{CVR}$ dynamic stability).
@@
A matching $m$ for $E^{T}$ is $\mathrm{CVR}$ dynamically stable if it is a $\Phi_{\hat{\mathrm{DS}}}$-solution for $E^{T}$. We denote by $\hat{\mathrm{DS}}_{T}(E^{T})$ the set of such solutions.
@@
The proof of 5.1 is in Appendix B. The proof proceeds by induction on the length of the economy, first showing the set of conjectures is nonempty and then showing they satisfy consistency; Theorem 4.1 then implies the set of $\mathrm{CVR}$ dynamically stable matchings is nonempty. Thus, the key step is showing the conjectures are nonempty: Because of the fixed-point character of the conjectures in period $t$, we build a recursion whose iteration delivers the set $\varphi_{\hat{\mathrm{DS}},t}$ as a limit.
@@
$\mathrm{CVR}$ dynamic stability is a refinement of dynamic stability, which is the stability notion in Doval (2022). In particular, if we denote by $\varphi_{\mathrm{DS}}$ the conjectures corresponding to dynamic stability, one can show that for all $t$, all $m^{t-1}$, and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, the conjectures $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k)\subseteq\varphi_{\mathrm{DS},t}(m^{t-1},k)$, which in turn implies the set of $\mathrm{CVR}$ dynamically stable matchings is a subset of those that are dynamically stable.
To understand the difference between the solution concepts, comparing the conjectures in Equation 5.3 with those in Doval (2022) is instructive. Whereas both solution concepts require agents to conjecture that the continuation matching conforms with the solution concept in the induced continuation economy (i.e., requirement (ii) in Equation 5.3), the conjectures in Doval (2022) impose a weaker requirement on the period-$t$ matching, namely, that among the agents who exit the economy in period $t$, the period-$t$ matching is individually rational and no blocking pairs exist.Footnote 17Footnote 17Footnote 17Formally, let $\varphi_{k}^{IR}$ denote the set of matchings in $\pazocal{M}_{T}(m^{t-1},k)$ that are individually rational for $k$. Then, in dynamic stability, matching $\overline{m}\in\pazocal{M}_{T}(m^{t-1},k)$ is a conjecture for $k$ if (i) $\overline{m}_{t}$ is stable among those who match relative to continuation values $\varphi_{\cdot}^{IR}$ and (ii) the continuation matching is dynamically stable. In other words, under dynamic stability, an agent $k$ conjectures that other contemporaneous agents are willing to exit in period $t$ with partners that dominate being unmatched through period $T$, even if they may not dominate the worst conjectured matching under $\varphi_{\mathrm{DS},t}(m^{t-1},\cdot)$. At the same time, when considering whether to block the proposed matching in period $t$, the agents available to match in period $t$ compare their payoff under the proposed matching with their worst conjectured matching under $\varphi_{\mathrm{DS},t}(m^{t-1},\cdot)$. That is, in dynamic stability, agents make their decisions to be available to match in period $t$ relative to their conjectures, but these conjectures do not take into account that other contemporaneous agents use their own conjectures when evaluating whether to match in period $t$. $\mathrm{CVR}$ dynamic stability eliminates this gap by requiring the period-$t$ agents to internalize that other contemporaneous agents’ continuation values are exactly those prescribed by their conjectures.
We conclude this section illustrating that $\mathrm{CVR}$ dynamic stability is a strict refinement of dynamic stability:
@@
The matching on the left, $m^{L}$, is a dynamically stable matching, with $m^{C}$ describing the conjectured matching that dissuades $a_{2}$ from waiting to be matched in $t=2$. Importantly, under $m^{C}$, $a_{1}$ and $b_{1}$ match in period $1$. However, we can easily see $a_{1}$ can always guarantee matching with $b_{3}$ (and similarly, $b_{1}$ with $a_{3}$) by remaining unmatched in $t=1$. In other words, $m^{C}$ cannot be a conjecture for $a_{2}$ under $\varphi_{\hat{\mathrm{DS}}}$, and hence, $m^{L}$ is not $\mathrm{CVR}$ dynamically stable. Instead, $m^{R}$ is $\mathrm{CVR}$ dynamically stable (and hence, dynamically stable).